# Spatiotemporal Semantic V2X Framework for Cooperative Collision Prediction

### v1

@@ Abstract
Intelligent Transportation Systems (ITS) demand real-time collision prediction to ensure road safety and reduce accident severity. Conventional approaches rely on transmitting raw video or high-dimensional sensory data from roadside units (RSUs) to vehicles, which is impractical under vehicular communication bandwidth and latency constraints. In this work, we propose a semantic V2X framework in which RSU-mounted cameras generate spatiotemporal semantic embeddings of future frames using the Video Joint Embedding Predictive Architecture (V-JEPA). To evaluate the system, we construct a digital twin of an urban traffic environment enabling the generation of diverse traffic scenarios with both safe and collision events. These embeddings of the future frame, extracted from V-JEPA, capture task-relevant traffic dynamics and are transmitted via V2X links to vehicles, where a lightweight attentive probe and classifier decode them to predict imminent collisions. By transmitting only semantic embeddings instead of raw frames, the proposed system significantly reduces communication overhead while maintaining predictive accuracy. Experimental results demonstrate that the framework with an appropriate processing method achieves a 10% F1-score improvement for collision prediction while reducing transmission requirements by four orders of magnitude compared to raw video. This validates the potential of semantic V2X communication to enable cooperative, real-time collision prediction in ITS.
@@ I Introduction
Traffic collisions remain one of the most severe challenges in Intelligent Transportation Systems (ITS), often leading to congestion, property damage, injuries, and fatalities. Proactive collision prediction is essential to mitigate these risks by enabling early warnings and preventive actions [9]. Traditional approaches that rely on transmitting raw video [16] or high-dimensional sensory data from roadside units (RSUs) to vehicles are not feasible under V2X constraints, where both bandwidth and latency are limited. This calls for communication-efficient, task-oriented methods that transmit only the most relevant information to support predictive safety applications.
@@
Building on this foundation, we propose a semantic V2X collision prediction framework. RSUs equipped with cameras employ the Video Joint Embedding Predictive Architecture (V-JEPA) as the semantic encoder [3]. V-JEPA does not merely summarize observed frames; it predicts the embeddings of future frames, capturing the evolving motion patterns and vehicle interactions that lead to collisions. These predicted embeddings are transmitted to vehicles via V2X links, where an attentive probe and classifier act as semantic decoders, refining the embeddings to predict whether a collision is imminent. By focusing on predictive embeddings, the system moves beyond detection toward true anticipation of dangerous events. The main contributions of this work are summarized as follows:
  • We construct a high-quality video dataset using a digital twin of an urban traffic environment (QLabs). The dataset includes diverse scenarios and is labeled into two categories, collision events and safe driving, enabling systematic evaluation of predictive collision avoidance.
  • We investigate multiple post-processing techniques (heatmaps, binary masking, and hybrid methods) to enhance task-relevant features in traffic videos, and analyze their impact on the representation quality and predictive performance of the V-JEPA framework.
  • We employ V-JEPA as the semantic encoder at RSUs to extract low-dimensional spatiotemporal embeddings of future frame from video clips. These embeddings serve as semantic messages transmitted via V2X links, and are decoded by an attentive probe with cross-attention at the vehicle side to focus on high-risk regions and motion patterns for accurate collision prediction.
This work demonstrates that combining digital twins, semantic communication, and spatiotemporal embeddings can enable predictive collision avoidance in vehicular networks, achieving both high accuracy and significant communication efficiency.
@@ II Related Works
Collision prediction and avoidance have been studied from multiple perspectives. The authors in [8] propose a deep learning-based approach using convolutional neural networks (CNNs) to analyze surveillance video for traffic congestion detection and accident monitoring, aiming to overcome the limitations of manual observation and traditional machine learning models. Similarly, [2] employs ensemble learning methods such as Random Forest, XGBoost, L-GBM, and CatBoost combined with Shapley Additive Explanations (SHAP) to predict accident severity and identify contributing factors, providing interpretable insights into road safety analysis. These works, however, largely focus on post-hoc analysis of accidents rather than proactive prediction of imminent collisions.
@@
YOLO-based detection frameworks are also gaining traction. In [10], YOLOv8 and YOLO11x models were trained on 15,000 traffic images for real-time accident and speeding-vehicle detection, introducing persistent multi-frame verification and human-in-the-loop validation. For proactive collision avoidance, [5] integrates CARLA simulations with YOLOv7-based object detection to anticipate collisions in connected and autonomous vehicle scenarios. Similarly, [17] enhances small-object detection and near-miss prediction by combining YOLOv7 with instance segmentation, Inverse Perspective Mapping (IPM), StrongSORT tracking, and Kalman filtering. These YOLO-based methods are effective in object-level detection and trajectory forecasting but still require transmitting large amounts of visual data for centralized processing.
In contrast, this work proposes a V2X SemComm framework where RSUs generate predictive spatiotemporal embeddings using V-JEPA. Unlike CNN- or YOLO-based methods that detect current events, V-JEPA forecasts embeddings of future frames, anticipating motion patterns that may lead to collisions. Instead of transmitting raw video, embeddings are sent as semantic messages to vehicles, which decode them with a lightweight attentive probe and classifier. This design enables real-time, bandwidth-efficient collision prediction, and, to our knowledge, is the first to integrate digital twin–based data generation, predictive embeddings, and semantic V2X communication.
@@ III Methodology
This work develops a semantic V2X framework in which RSUs equipped with surveillance cameras and V-JEPA act as semantic encoders, and vehicles act as semantic decoders for predictive collision avoidance. Unlike image-based approaches, which provide only static snapshots, video data captures both spatial and temporal dynamics, enabling richer modeling of vehicle interactions, motion patterns, and speed estimation. V-JEPA is employed at the RSU to shift predictive modeling from pixel-level forecasting to the representation space. Instead of reconstructing raw frames, V-JEPA generates compact spatiotemporal embeddings of future frames, thereby anticipating anomalous motion patterns that precede collisions. These predicted embeddings are transmitted as semantic messages over V2X links to vehicles, where an attentive probe and classifier decode them to determine whether a collision is imminent or the driving condition is safe. By transmitting embeddings rather than raw video, the system reduces both communication overhead and computational workload while preserving predictive accuracy.
@@ III-A1 Video Generation using QLabs
To obtain a realistic and diverse dataset, the QLabs platform is employed as a digital twin of an urban traffic environment [15]. QLabs models intersections, roundabouts, two-way roads, pedestrians, traffic lights, and autonomous vehicles, enabling flexible and reproducible scenario creation. The Qlabs environment is programmed to generate both safe driving and collision events, observed from four RSU-mounted surveillance cameras covering different regions of the environment. Each frame is sampled every 50 ms, producing 20 FPS (Frames per second) video clips of varying lengths depending on scenario duration. This allows for reproducible traffic scene generation and controlled experimentation for predictive safety analysis. Figure1 depicts the process of video-dataset generation.
@@ III-B Semantic Encoder: V-JEPA at RSU
The V-JEPA model [3] is deployed at RSUs to generate predictive semantic embeddings. Unlike conventional predictive models, V-JEPA forecasts embeddings of masked future frame regions rather than reconstructing pixels, thereby focusing on task-relevant semantics for downstream prediction.
@@ III-B1 V-JEPA Pretraining with Masked Modeling
As shown in Figure3, V-JEPA is pretrained in a self-supervised fashion using masked prediction. A video clip is divided into non-overlapping spatiotemporal patches. The target encoder processes the full unmasked clip, while the context encoder encodes a masked version. A predictor network generates embeddings for the masked tokens, which are trained against target encoder embeddings using an $L_{1}$ loss. This allows the model to learn predictive spatiotemporal representations in the embedding space.
Figure 3: V-JEPA Pretraining & Fine-Tuning for Collision Prediction and V2X Semantic Communication
@@ III-B2 Fine-Tuning with Attentive Probe
For downstream collision prediction, the pretrained V-JEPA encoder takes as input a video clip of dimension $N\times W\times H$, where $N$ is the number of frames, and $W$ and $H$ denote frame width and height. The clip is flattened into a sequence of non-overlapping spatiotemporal patches, producing token embeddings of shape $N\times P\times D$, where $P$ is the number of patches per frame and $D$ is the embedding dimension. These frozen encoder parameters preserve the rich self-supervised features learned during pretraining. On top of the encoder outputs, an attentive probe applies a query-driven cross-attention mechanism to pool the tokens into a single compact vector $1\times D$, selectively aggregating information from the most relevant spatial and temporal regions. This pooled embedding is then passed to a lightweight classifier to infer one of two states: collision or safe driving. By transforming raw video data ($N\times W\times H$) into a compact embedding $1\times D$, the framework enables efficient semantic transmission while maintaining predictive accuracy for V2X communication.
@@ III-C Semantic Decoder at Vehicle
At the vehicle side, the received embeddings are directly processed by the semantic decoder, which consists of an attentive probe and a lightweight classifier. Vehicles receive embeddings from RSUs, refine them through query-driven cross-attention, concentrating on the most critical spatial and temporal cues that may indicate anomalous motion or collision risk. The resulting compact representation is then fed into a linear classifier to produce a binary decision: collision or safe driving.
@@ III-D Transmission Efficiency in V2X Semantic Communication
A major advantage of the proposed framework lies in its ability to drastically reduce transmission payloads compared to conventional raw video communication. Let each video clip contain $N$ frames of original spatial size $H_{o}\times W_{o}$ with 3 RGB channels, where $(H_{o},W_{o})$ denote the preprocessed frame dimensions. The total raw video payload can be expressed as:
@@
In contrast, the semantic embeddings generated by the V-JEPA encoder at the RSU are significantly more compact. Each clip produces embeddings of size $N\times P\times D$, where $P$ is the number of spatiotemporal patches per frame and $D$ is the embedding dimension. The corresponding payload is:
$\small S_{\text{sem}}=N\cdot P\cdot D\cdot b\quad\text{bytes},$ (2)
where $b$ denotes the number of bytes per element ($b=4$ for FP32, $b=2$ for FP16, $b=1$ for INT8). The semantic compression ratio is then: $\mathcal{R}={S_{\text{raw}}}/{S_{\text{sem}}}$. Assuming $N=64$, $H_{o}=W_{o}=2048$, and a post-processed input of $384\times 384$ with patch size $16\times 16$ ($P=\tfrac{384}{16}\times\tfrac{384}{16}=576$) and embedding dimension $D=1280$, the corresponding payloads are:
  • Raw video (RGB): 0.81 GB per clip.
  • Embeddings (FP32): 0.74 MB ($\mathcal{R}\approx 1095\times$).
  • Embeddings (FP16): 0.37 MB ($\mathcal{R}\approx 2190\times$).
  • Embeddings (INT8): 0.18 MB ($\mathcal{R}\approx 4380\times$).
These results demonstrate that transmitting semantic embeddings instead of raw video frames achieves up to four orders of magnitude reduction in communication payload. This enables low-latency, bandwidth-efficient semantic message exchange between RSUs and vehicles, ensuring timely and reliable collision prediction in V2X networks.
@@ III-E2 Attentive Probe and Classifier Cost
where $C$ denotes the number of output classes (e.g., $C=2$ for collision and safe driving). However, in practice, the encoder output is frozen and the probe operates with a single query vector rather than performing full self-attention over all $L$ tokens. Consequently, the cost associated with key and value projections ($\mathcal{O}(LD^{2})$) is computed once during encoding and not repeated during probing. The probe therefore performs only lightweight attention ($\mathcal{O}(LD)$) and projection operations ($\mathcal{O}(D^{2})$), resulting in an overall complexity that scales as:
$\small\mathcal{O}(D^{2}+DC),\small$ (5)
@@ III-E3 End-to-End Computational Cost
If $V$ different spatiotemporal views (e.g., crops or augmentations) are processed per clip, the total cost scales linearly with $V$. Given a device with computational throughput $\phi$ (in FLOPs/s) and per-view input/output latency $t_{\text{I/O}}$, the total inference time can be approximated as:
$\small t_{\text{infer}}={F_{\text{total}}}/{\phi}+V\,t_{\text{I/O}}.$ (7)
Hence, the encoder dominates the total computational cost, while the attentive probe and classifier introduce only minor overhead ($\mathcal{O}(D^{2}+DC)$). This computational asymmetry justifies deploying the encoder at the RSU (with higher compute resources) and the lightweight decoder on the vehicle side.
IV Performance Evaluations
@@ IV-A Evaluation Setup and Parameters
In the simulation setup, as mentioned in Section III-A1, four different cameras are used to collect consecutive frames at 50 ms intervals, which are then used to generate video clips at 20 FPS. In total, 500 traffic scenario clips have been collected, with 115 including collisions and 385 consisting of safe driving [^1] Urban Mobility Research Video Dataset (Generated with Quanser Interactive Lab), Available at: https://tinyurl.com/Uottawa-vjepa. Each frame of the videos has a 2048x2048 dimension, and the durations vary depending on the simulation runtime. The simulation automatically finalizes itself once a collision occurs. Each video’s total lengths are adjusted to a maximum of 64 frames by removing from the beginning to make it adaptable for V-JEPA training parameters. For various post-processing methods, explained in Section III-A2, a previously trained YOLOv11 model is employed from [11].
For the V-JEPA model, a ViT-Huge (ViT-H) based encoder model, which is trained using masked modeling self-supervised learning, as explained in Section III-B1, is employed directly from its repository [3], having been pretrained with VideoMix2M beforehand. ViT-H requires 384x384 video frames, so each clip, before forwarding to the V-JEPA encoder, is resized to the proper shape. Moreover, the patch size in its training is a 16x16 pixel block, and the embedding output dimension of ViT-H is a 1280-sized vector, which is used in SemComm. For fine-tuning to a specific task, as explained in Section III-B2, the K400 attentive probe is utilized. After video clips are forwarded to the frozen encoder of V-JEPA, the K400 attentive probe is trained with 40 epochs and an 8-batch size. The output layer of the attentive probe consists of two nodes: Safe Driving and Collision Probability. During the training, the AdamW optimizer is utilized with 0.001 initial learning rate of 0.001. All training parameters can be seen in Table I.
TABLE I: Performance Evaluation Setup

### v2

@@ Abstract
Intelligent Transportation Systems (ITS) demand real-time collision prediction to ensure road safety and reduce accident severity. Conventional approaches rely on transmitting raw video or high-dimensional sensory data from roadside units (RSUs) to vehicles, which is impractical under vehicular communication bandwidth and latency constraints. In this work, we propose a semantic V2X framework in which RSU-mounted cameras generate spatiotemporal semantic embeddings of future frames using the Video Joint Embedding Predictive Architecture (V-JEPA). Afterwards, These embedding future frames are utilized to estimate the likelihood collision for alert the prevention system at vehicle side. To evaluate the system, we construct a digital twin of an urban traffic environment enabling the generation of diverse traffic scenarios with both safe and collision events. The embeddings of the future frame, extracted from V-JEPA, capture task-relevant traffic dynamics and are transmitted via V2X links to vehicles, where a lightweight classifier decode them to predict imminent collisions. By transmitting only semantic embeddings instead of raw frames, the proposed system significantly reduces communication payload while maintaining predictive accuracy. Experimental results demonstrate that the proposed framework with an appropriate processing method achieves 92% accuracy and 8% F1-score improvement for collision prediction. It also reduces transmission payload by five orders of magnitude compared to raw video fulfilling the latency constraints for safe driving. This validates the potential of semantic V2X communication to enable cooperative, real-time collision prediction in ITS.
@@ I Introduction
Achieving proactive collision prediction is essential for enhancing road safety through early warnings and coordinated preventive actions in Intelligent Transportation Systems (ITS) [1]. However, conventional approaches that transmit raw video streams [2] or high-dimensional sensor data from roadside units (RSUs) to vehicles are impractical under vehicular communication (V2X) constraints due to limited bandwidth and stringent latency requirements. These limitations highlight the need for communication-efficient, task-oriented frameworks that convey only semantically relevant information to support predictive safety functions in connected and autonomous vehicles.
@@
Building on this foundation, we propose a novel semantic V2X collision prediction framework which departs from conventional perception-centric V2X approaches by performing future-oriented semantic prediction. RSUs equipped with cameras employ the Video Joint Embedding Predictive Architecture (V-JEPA) as the semantic encoder [4]. V-JEPA does not merely summarize observed frames; it predicts the embeddings of future frames, capturing the evolving motion patterns and vehicle interactions that lead to collisions. These predicted embeddings are transmitted to vehicles via V2X links, where a classifier acts as a semantic decoder, refining the embeddings to predict whether a collision is imminent. By focusing on predictive embeddings, the system moves beyond detection toward true anticipation of dangerous events. The main contributions of this work are summarized as follows:
  • A high-quality video dataset is generated using a QLabs digital twin, featuring diverse urban scenarios labeled as collision or safe, supporting systematic evaluation of predictive collision avoidance.
  • We evaluate post-processing methods such as heatmaps, binary masks, and hybrids to enhance task-relevant features and assess their effect on V-JEPA’s representation quality and predictive accuracy.
  • V-JEPA is employed to encode future-frame embeddings at RSUs, which are transmitted via V2X and decoded by a lightweight classifier in vehicles to identify high-risk regions for collision prediction.
This work demonstrates that combining digital twins, SemComm, and spatiotemporal embeddings can enable predictive collision avoidance in vehicular networks, achieving 92% accuracy and 8% F1-score improvement in collision prediction with five orders of magnitude reduction in transmission (Tx) payload, supporting real-time latency constraint V2X communication.
@@ II Related Works
Collision prediction and avoidance have been explored from various perspectives. In [5], a deep learning model based on convolutional neural networks (CNNs) analyzes surveillance video for traffic congestion and accident detection, addressing the limits of manual observation and conventional learning methods. Similarly, the work in [6] uses ensemble models (Random Forest, XGBoost, L-GBM, CatBoost) with SHAP-based interpretability to assess accident severity and contributing factors. However, these studies mainly emphasize post-event analysis rather than proactive prediction of collisions.
@@
YOLO-based detection frameworks have gained attention for traffic analysis. In [11], YOLOv8 and YOLO11x models are trained on 15,000 traffic images for real-time accident and speeding-vehicle detection. For proactive collision avoidance, the work in [12] integrates CARLA simulations with YOLOv7-based object detection, while the study in [13] enhances small-object detection and near-miss prediction through instance segmentation, IPM, and tracking modules. Although effective for object-level detection, these methods still rely on transmitting large volumes of visual data for centralized processing.
In contrast, this work introduces a V2X semantic communication framework where RSUs use V-JEPA to generate predictive spatiotemporal embeddings of future frames, enabling anticipation of collision-prone motion patterns. Rather than transmitting raw video, semantic embeddings are shared with vehicles for decoding via a lightweight classifier, achieving real-time and bandwidth-efficient collision prediction.
@@ III Methodology
This work develops a semantic V2X framework in which RSUs equipped with surveillance cameras and V-JEPA act as semantic encoders, and vehicles act as semantic decoders for predictive collision avoidance. Unlike image-based approaches, which provide only static snapshots, video data captures both spatial and temporal dynamics, enabling richer modeling of vehicle interactions, motion patterns, and speed estimation. V-JEPA is employed at the RSU to shift predictive modeling from pixel-level forecasting to the representation space. Instead of reconstructing raw frames, V-JEPA generates compact spatiotemporal embeddings of future frames, thereby anticipating anomalous motion patterns that precede collisions. These predicted embeddings are transmitted as semantic messages over V2X links to vehicles, where a classifier decode them to determine whether a collision is imminent or the driving condition is safe. By transmitting embeddings rather than raw video, the system reduces both communication payload and computational workload while preserving predictive accuracy.
@@ III-A1 Video Generation using QLabs
To obtain a realistic and diverse dataset, the QLabs platform is employed as a digital twin of an urban traffic environment [14]. QLabs models intersections, roundabouts, two-way roads, pedestrians, traffic lights, and autonomous vehicles, enabling flexible and reproducible scenario creation. The Qlabs environment is programmed to generate both safe driving and collision events, observed from four RSU-mounted surveillance cameras covering different regions of the environment. Simulation provides 4 different type of regions: 4-way junction, 3-way junction, side roads and roundabout. There are total 500 video clips and each frame is sampled every 50 ms, producing 20 FPS (Frames per second) of varying lengths depending on scenario duration. This allows for reproducible traffic scene generation and controlled experimentation for predictive safety analysis. Figure1 depicts the process of video-dataset generation.
@@ III-B Semantic Encoder: V-JEPA at RSU
The V-JEPA model [4] is deployed at RSUs to generate predictive semantic embeddings. Unlike conventional predictive models, V-JEPA forecasts embeddings of masked future frame rather than reconstructing pixels, thereby focusing on task-relevant semantics for downstream prediction.
@@ III-B1 V-JEPA Pretraining with Masked Modeling
As shown in Figure 3, V-JEPA is pretrained in a self-supervised fashion using masked prediction. A video clip is divided into non-overlapping spatiotemporal patches, where each frame of spatial size $H\times W$ is partitioned into $P=\tfrac{H}{p}\times\tfrac{W}{p}$ patches based on a patch size of $p\times p$. The target encoder processes the full unmasked clip, while the context encoder encodes a masked version. Each patch is projected into a $D$-dimensional embedding vector, resulting in a per-frame embedding tensor of size $P\times D$. A predictor network then generates embeddings for the masked tokens, which are trained against the corresponding target encoder embeddings using an $L_{1}$ loss. This enables the model to learn predictive spatiotemporal representations directly in the embedding space rather than at the pixel level.
Figure 3: V-JEPA Pretraining and Fine-Tuning for Collision Prediction
@@ III-B2 Fine-Tuning with Attentive Probe
For downstream collision prediction, the pretrained V-JEPA encoder takes as input a video clip of dimension $N\times W\times H$, where $N$ is the number of frames in the clip. The clip is flattened into a sequence of non-overlapping spatiotemporal patches, producing token embeddings of shape $N\times P\times D$, where $P$ is the number of patches per frame and $D$ is the embedding dimension. These frozen encoder parameters preserve the rich self-supervised features learned during pretraining. On top of the encoder outputs, The attentive probe employs a query-driven cross-attention mechanism, where the query vector acts as a learnable global token that attends to the most informative spatial and temporal regions in the video. Formally, let the encoder output token embeddings be $\mathbf{Z}\in\mathbb{R}^{(N\cdot P)\times D}$. The attentive probe introduces a learnable query matrix $\mathbf{Q}\in\mathbb{R}^{Q\times D}$, with $Q=1$ in this setup, to perform cross-attention as follows:
$\small\mathbf{z}_{\text{attn}}=\text{Attn}(\mathbf{Q},\mathbf{Z})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{Z}^{\top}}{\sqrt{D}}\right)\mathbf{Z},$ (1)
where $\mathbf{Q}$ serves as the query, and $\mathbf{Z}$ provides both keys and values. The resulting weighted aggregation yields a single compact embedding $\mathbf{z}_{\text{attn}}\in\mathbb{R}^{1\times D}$, which captures the most salient spatiotemporal dynamics relevant to collision prediction. This pooled embedding is then passed to a lightweight classifier to infer one of two states: collision or safe driving. By transforming raw video data ($N\times W\times H$) into a compact embedding $1\times D$, the framework enables efficient semantic Tx while maintaining predictive accuracy for V2X communication.
@@ III-C Semantic Decoder at Vehicle
At the vehicle side, the received embeddings are directly processed by the semantic decoder, which consists of a lightweight classifier. Vehicles receive embeddings from RSUs, that are refined through query-driven cross-attention, concentrating on the most critical spatial and temporal cues that may indicate anomalous motion or collision risk. The resulting compact representation is then fed into a linear classifier to produce a binary decision: collision or safe driving.
@@ III-D Transmission Efficiency in V2X Semantic Communication
A major advantage of the proposed framework lies in its ability to drastically reduce Tx payloads compared to conventional raw video communication. Let each video clip contain $N$ frames of original spatial size $H_{o}\times W_{o}$ with 3 RGB channels, where $(H_{o},W_{o})$ denote the preprocessed frame dimensions. The total raw video payload can be expressed as:
@@
In contrast, the semantic embeddings generated by the V-JEPA encoder at the RSU are significantly more compact. Each clip produces embeddings of size $N\times P\times D$, where $P$ is the number of spatiotemporal patches per frame and $D$ is the embedding dimension. Then, an attentive probe aggregates these tokens into a single compact representation of size $1\times D$ using a query-driven cross-attention mechanism. This process selectively emphasizes the most informative spatial and temporal regions, resulting in a lightweight semantic vector suitable for V2X transmission. The payload corresponding to the transmitted semantic embedding is therefore:
$\small S_{\text{sem}}=1\times D\times b\quad\text{bytes},$ (3)
where $b$ denotes the number of bytes per element ($b=4$ for FP32, $b=2$ for FP16, $b=1$ for INT8). The semantic compression ratio can be defined as $\mathcal{R}={S_{\text{raw}}}/{S_{\text{sem}}}.$
@@ III-E2 Attentive Probe and Classifier Cost
where $C$ denotes the number of output classes (e.g., $C=2$ for collision and safe driving). However, in practice, the encoder output is frozen and the probe operates with a single query vector rather than performing full self-attention over all $L$ tokens to capture a unified global representation. Consequently, the cost associated with key and value projections ($\mathcal{O}(LD^{2})$) is computed once during encoding and not repeated during probing. The probe therefore performs only lightweight attention ($\mathcal{O}(LD)$) and projection operations ($\mathcal{O}(D^{2})$), resulting in an overall complexity that scales as:
$\mathcal{O}(D^{2}+DC),\small$ (6)
@@ III-E3 End-to-End Computational Cost
If $V$ different spatiotemporal views (e.g., crops or augmentations) are processed per clip, the total cost scales linearly with $V$. Given a device with computational throughput $\phi$ (in FLOPs/s) and per-view input/output latency $t_{\text{I/O}}$, the total inference time can be approximated as $t_{\text{infer}}={F_{\text{total}}}/{\phi}+V\,t_{\text{I/O}}.$ Hence, the encoder dominates the total computational cost, while the attentive probe and classifier introduce only minor overhead ($\mathcal{O}(D^{2}+DC)$). This computational asymmetry justifies deploying the encoder at the RSU (with higher compute resources) and the lightweight decoder on the vehicle side.
IV Performance Evaluation
@@ IV-A Evaluation Setup and Parameters
As mentioned in Section III-A1, four different cameras are used in the QLabs platform to collect consecutive frames at 50 ms intervals, which are then used to generate video clips at 20 FPS. In total, 500 traffic scenario clips have been collected, with 115 including collisions and 385 consisting of safe driving [^1] Urban Mobility Research Video Dataset, Available at: https://tinyurl.com/Uottawa-vjepa. Generated with Quanser Interactive Lab: (https://www.quanser.com/products/quanser-interactive-labs/). The clips’ durations vary depending on the simulation runtime. The simulation automatically finalizes itself once a collision occurs. The long clips’ lengths are adjusted to a maximum of 64 frames by removing from the beginning to make it adaptable for V-JEPA training parameters. For various post-processing methods, explained in Section III-A2, a previously trained YOLOv11 model is employed from [16].
For the V-JEPA model, a ViT-Huge (ViT-H) encoder pretrained with VideoMix2M via masked modeling self-supervised learning [4] is employed, as described in Section III-B1. Each video clip is resized to 384×384 before encoding, with a patch size of 16×16 and an embedding dimension of 1280 used in the SemComm framework. For task-specific fine-tuning (Section III-B2), the K400 attentive probe is trained for 40 epochs with a batch size of 8, using the AdamW optimizer and an initial learning rate of 0.001. The output layer includes two nodes, namely the Safe Driving and Collision Probability, and the training parameters are summarized in Table I.
TABLE I: Training Setup and Parameters
@@
Modulation Schemes/SNR BPSK/12dB, QAM-16/22dB [17]
Encoding Format INT8, FP16, FP32
Bandwidth 20 MHz